A rigid tank contains 1.50 moles of an ideal gad. Determine the number of loes of gas that must be withdrawn from the tank to lower the pressure of the fas from 25.0 atm to 5.00 atm. Assume the volume of the tank and the temperature of the gas remain constant during this operation.
1
Expert's answer
2021-02-19T10:32:35-0500

"p_1V_1=\nu_1RT_1,\\p_2V_2=\nu_2RT_2,"

"V_1=V_2,~T_1=T_2,"

"\frac{p_1}{p_2}=\frac{\nu_1}{\nu_2},"

"\nu_2=\nu_1\frac{p_2}{p_1},"

"\Delta \nu=\nu_1-\nu_2=\nu_1\frac{p_1-p_2}{p_1},"

"\Delta\nu=1.2~\text{moles.}"